### v1

# Compositional Coordination for Multi‑Robot Teams with Large Language Models

Abstract
Multi-robot coordination has traditionally relied on a task-specific and expert-driven pipeline, where natural language mission descriptions are manually translated by domain experts into mathematical formulation, algorithm design, and executable code. This conventional process is labor-intensive, inaccessible to non-experts, and inflexible to changes in mission requirements. Here, we propose LAN2CB (Language to Collective Behavior), a novel framework that leverages large language models (LLMs) to streamline and generalize the multi-robot coordination pipeline. LAN2CB directly converts natural language mission descriptions into executable Python code for multi-robot systems through two key components: (1) Mission Decomposition for Task Representation, which parses the mission into a task graph with dependencies, and (2) Code Generation, which uses the task graph and a structured knowledge base to generate deployable robot control code. We further introduce a dataset of natural language mission specifications to support development and benchmarking. Experimental results in both simulation and real-world settings show that LAN2CB enables effective and flexible multi-robot coordination from natural language, significantly reducing the need for manual engineering while supporting generalization across mission types. Website: https://sites.google.com/view/lan2cb
I Introduction
Problem: Multi-robot coordination for complex collective behaviors, including coverage, formation, foraging, and exploration, has been extensively studied for decades [1, 2, 3, 4, 5, 6, 7]. However, most prior coordination methods target a specific and narrowly defined coordination task. This highlights the need for a general and unified framework that can address a wide range of coordination problems in a systematic way. The conventional multi-robot development pipeline starts with a natural language (NL) problem description, enabling both expert and non-expert users to specify objectives and constraints. An expert translates this description into a mathematical problem (typically an optimization problem). This inevitably involves making some assumptions. Following this, a domain expert designs algorithms to solve the formulated problem, which are subsequently implemented and validated on robotic hardware. However, this pipeline has three limitations: It is labor-intensive, inaccessible to non-experts, and inflexible. The first two limitations are apparent due to the significant use of skilled experts, despite the repetitive nature of problem formulation, algorithm design, and code implementation. The third becomes apparent when small changes to the problem statement or new requirements emerge, and the entire pipeline has to be manually redesigned.
Claim: We posit that recent advances in Large Language Models (LLMs) can be leveraged to alleviate these three limitations. Specifically, the use of such models can significantly reduce human expert effort and enhance the flexibility of multi-robot coordination systems [8, 9, 10, 11, 12].
Background: LLMs are trained on massive amounts of text data to understand and generate human-like language, allowing them to interpret natural language prompts, perform contextual reasoning, and produce coherent responses or actions in both visual and linguistic tasks[13, 14, 15, 16, 17, 18, 19]. In robotics, LLMs are increasingly being used to improve high-level decision-making, task planning, and human-robot interaction [20, 21, 22, 23, 24, 25, 26, 27, 28]. Applications include interpreting natural language instructions, generating structured plans, assisting in code generation for robot control, and enabling dialogue-based coordination among multi-robot systems. The key advantages of LLMs in robotics are their flexibility, adaptability to new tasks without extensive retraining, and ability to bridge the gap between human intent and machine execution, significantly reducing the need for manual programming and domain-specific engineering.
Figure 1: LAN2CB is an LLM-assisted multi-robot framework that converts mission descriptions into structured representations (e.g., behavior trees), assigns roles and priorities to each robot, and automatically generates executable code to accomplish complex missions. A mission change does not require human intervention since new code is generated automatically. NL means natural language.   user- or expert-provided content.   LLM-generated mission representations.   LLM-generated code.
Contributions: 1. We propose, implement, and thoroughly evaluate an LLM-based framework, Language to Collective Behavior (LAN2CB), which addresses the limitations of the traditional pipeline and directly turns natural language into executable Python code for robot teams. LAN2CB consists of two main modules (Fig. 1). The first module is Mission Decomposition for Task Representation, which takes the mission specified in natural language as input, identifies the necessary tasks to be completed, and analyzes the dependency between tasks. The output of the first module, all ready-to-run nodes from a task graph, is given to the second module Code Generation, together with a preconstructed knowledge base, to generate executable code for robots. 2. We design a dataset of natural language mission specifications for multi-robot coordination. 3. Our experiments, both in simulation and hardware, demonstrate that LAN2CB produces effective multi-robot coordination from natural language. They also show that LAN2CB is flexible and extendable, obviating the need to design from scratch whenever the mission changes and providing a standard format to incorporate existing knowledge into its knowledge base.
II Related Work
II-A Multi-Robot Coordination
Multi-robot systems can solve large tasks efficiently through cooperation, coordination, or collaboration [29]. Prior work has shown that these approaches enable robots to contribute effectively to overall team performance, either through global control strategies [30] or by relying on local observations and decentralized decision-making [31, 32]. As tasks become more complex, task-level coordination has received growing attention, where high-level goals are decomposed into interdependent sub-tasks that are allocated and executed collaboratively by the team [33, 34, 35]. For example, Messing et al. [36] introduced a four-layer structure covering task planning, allocation, scheduling, and motion planning to simplify decision-making. More recent work applied learning-based methods to multi-robot coordination by leveraging neural architectures such as graph attention networks for scheduling [37] and heterogeneous policy networks for communication and collaboration in diverse robot teams [38]. However, most existing approaches remain limited to specific tasks and lack the generalization to handle more dynamic objectives, understand natural language, or perform higher-level reasoning abilities.
II-B Integrating LLMs in Robotics
II-B1 Single-Robot Integration
Recent work has explored integrating large language models (LLMs) with single robots to enable more intuitive human-robot interaction and high-level task planning. These approaches leverage the language understanding and reasoning capabilities of LLMs to interpret natural language commands and translate them into executable actions. Because natural language is inherently tied to contexts and semantics, LLMs are particularly effective in understanding tasks such as navigation, exploration, and instruction following. In these scenarios, LLMs can help robots ground language inputs into spatial or semantic representations that guide behavior in complex, partially known environments [23, 24, 22]. In addition, real-time residency [20] and anomaly detection [25] have been integrated into LLM-based robotic pipelines to improve robustness in dynamic and unforeseen situations.
II-B2 Team-Level Integration
Extending LLMs to multi-robot systems introduces challenges in coordination, communication, and scalability, with research still in the early stages exploring their use for flexible, robust multi-robot coordination [39]. Some works use LLMs to translate natural language into formal representations [40, 41, 42, 43], but they are limited to simple missions and cannot handle multi-robot teams or decomposable long-horizon tasks. Others propose prompt-based frameworks for multi-robot coordination using LLMs [27, 28]. For instance, Zhao et al. introduced RoCO [27], where robots coordinate via interactive dialogues. However, such frameworks face limitations with large robot teams (e.g., $\geq 10$) and complex language-based missions. Shyam et al. proposed SMART-LLM [28], an LLM-based task planner for multi-robot teams with task decomposition and allocation, but it lacks domain knowledge, limits generalization beyond in-context examples, and cannot handle motion-level language commands. Our work is closely related to GenSwarm [44], which translates natural language missions into executable Python code. While both aim to enable language-driven multi-robot coordination, our framework differs in two key ways. First, our modular design supports a broader range of coordination problems, including multi-team tasks with interdependencies, unlike GenSwarm’s focus on simple, single-team scenarios. Second, we incorporate a structured knowledge base to guide the LLM, enhancing extensibility, whereas GenSwarm relies solely on in-context learning, limiting generality.
Figure 2: A comprehensive workflow demonstration of LAN2CB.   user- or expert-provided content.   LLM-generated mission descriptions.   LLM-generated code.
III Problem Statement
Consider an environment with $N$ robots. Given natural language mission descriptions from users, our goal is to develop a system that can automatically control robots to accomplish these missions without further human intervention. We focus on two fundamental categories of missions: 1) goal-approaching missions, such as geometric formation or region visitation, and 2) herding missions, where robots must guide dynamic objects into specific regions or achieve a specific coverage percentage of an area. Mission diversity is introduced by systematically varying key dimensions, resulting in a broad spectrum of possible tasks:
  1. 1.
    Coordination Strategies: Missions may require robots to coordinate, form a range of geometric patterns (e.g., lines, circles, characters), visit assigned regions, track dynamic targets, or herd dynamic objects.
  2. 2.
    Motion Patterns: Robots may be required to execute various motion trajectories, such as moving in straight lines, zigzagging, or spiral paths.
  3. 3.
    Constraints: Additional requirements can be imposed, such as entering a region from a specific direction, maintaining speed limits, or adhering to prescribed durations for actions.
  4. 4.
    Trigger: Tasks can include trigger conditions, enabling actions to be terminated or modified in response to environmental events (e.g., if a robot comes within a certain distance, say 1 m, of a forbidden region, it switches to an alternative goal).
  5. 5.
    Atomic Task Finish Condition: Each atomic task has a clearly defined completion criterion, such as reaching a designated position within a user-specified tolerance or herding objects within a given range of the target region.
  6. 6.
    Mission Finish Condition: Mission completion is defined more broadly, either by the successful completion of all constituent atomic tasks or by satisfying higher-level objectives (e.g., achieving at least 60% coverage of a specified region).
IV Method
The LAN2CB workflow consists of the following stages, illustrated in Fig. 1:
  1. 1.
    Mission Analysis: Given natural language mission descriptions from users, LLMs autonomously decompose missions into atomic task specifications and analyze intertask dependencies. For each task, LLMs extract robot assignments, relevant constraints, triggers, finish conditions, contextual hints, and other contextual requirements. Tasks that are immediately executable are then identified and forwarded to the next stage.
  2. 2.
    Code Generation: Based on the mission description, the system first generates the mission-level completion criteria, which are produced once for the entire mission cycle. For each selected task, considering its constraints, trigger conditions, finish conditions, and hints, the system separately generates: (a) robot execution code, (b) code for evaluating trigger conditions, and (c) code for assessing task completion.
  3. 3.
    Execution: The generated code is deployed to the robots for execution. During run-time, if any trigger condition or task completion criterion is met, the workflow transitions to stage 4. If the overall mission completion condition is satisfied, it proceeds directly to stage 5.
  4. 4.
    Mission Progress Update: Upon receiving execution results, the system uses the LLM to update the status of each task, then returns to stage 1 for further analysis and potential replanning if necessary.
  5. 5.
    Mission Completion: Once all mission requirements are met, the workflow is terminated and the mission is formally marked as complete.
IV-A Mission Analysis
Mission analysis can be divided into two phases: mission understanding and ready-to-run action node selection. However, user-provided natural language mission descriptions often lack explicit definitions of all atomic tasks and their execution dependencies. Consequently, directly prompting LLMs for general mission understanding may result in incomplete or ambiguous task specifications and overlook critical execution dependencies. To address these challenges, we first standardized mission descriptions by using an expert-provided template, and then decomposed the mission understanding stage into three steps: 1) dependency analysis, 2) comprehensive mission analysis, and 3) ready-to-run action node(s) selection.
Dependency Analysis: To initiate dependency analysis, we prompt an LLM to extract all atomic tasks from the mission description, initially disregarding trigger conditions and any tasks that may be activated post-trigger. The LLM then analyzes the dependency relationships among these atomic tasks. If triggers are activated during execution, we re-invoke the LLM to update the dependency analysis accordingly.
Comprehensive Mission Analysis: Building on the results of dependency analysis, we combine these atomic tasks and their dependencies with the standardized mission description to prompt the LLM for comprehensive mission analysis. The mission is represented in a behavior tree structure, with the overall mission as the root and atomic tasks as action nodes. Non-leaf nodes of type Parallel or Sequence capture task execution topologies according to dependencies. For each node, we instruct the LLM to generate a unique index, identify the node type, task name, status, constraints, trigger conditions, finish conditions, hints, and children. For action nodes, the LLM also specifies the action type as well as the associated robots, objects, regions, and points. See Tab. I for detailed field descriptions of each BehaviorTreeNode.
TABLE I: Field Descriptions for BehaviorTreeNode.
Field (All Nodes) Description
idx Unique index of the node.
node_type Parallel |||| Sequence |||| Action.
task_name Concise task description of the node.
status Idle |||| Running |||| Success |||| Failure.
constraints Constraints such as max speed, time limit, regions to avoid, and etc.
trigger_condition Condition(s) that trigger node termination and activate new nodes.
finish_condition Condition(s) for determining task completion.
hints Hints or suggestions.
children List of child BehaviorTreeNode objects (empty for Action nodes).
Field (Action Nodes) Description
action_type visit_points |||| follow_targets |||| herd.
robot_ids List of robot IDs this action node will control.
object_ids List of object IDs related to the action node.
region_ids List of region IDs related to the action node.
points List of points related to the action node.
Ready-to-run Action Node(s) Selection: After constructing the complete behavior tree and integrating the dependency analysis results, we ask LLMs to identify action nodes that are immediately executable. Specifically, these are nodes for which all dependencies are satisfied and all preconditions are met in the current environment state. The ready-to-run action nodes are then selected and forwarded for downstream processing, code generation. This approach ensures that robot actions are initiated only when all prerequisite tasks and constraints have been resolved, thereby supporting safe and efficient mission execution. After selecting action nodes that are ready-to-run, we instruct LLMs to update the status of all nodes in the behavior tree, marking the selected nodes and their parent nodes as Running.
IV-B Code Generation
After mission analysis and the identification of ready-to-run action nodes, we prompt LLMs to generate Python code to control all robots in the environment. The code generation process consists of four key components, executed synchronously with the robots’ control loop (Fig. 1 and  2):
  1. 1.
    Robot execution code generation;
  2. 2.
    Trigger condition code generation;
  3. 3.
    Action node finish condition code generation;
  4. 4.
    Mission finish condition code generation.
Robot Execution Code Generation: Recent works have shown that LLMs can effectively assist code generation [45, 46, 47, 48]. However, to increase the success rate and avoid hallucination, we still need to provide enough prior information to LLMs instead of asking LLMs to generate all code from scratch [49]. To this end, we construct a meta behavior library for robot execution. To ensure flexibility and the capability to process multiple action nodes simultaneously, the meta-behavior library comprises three sub-libraries:
  • Goal Generation: Provides primitives for three fundamental behaviors: (a) visiting points, (b) following targets, and (c) herding.
  • Goal Allocation: Offers two strategies for mapping generated goals to robots: (a) a default mapping that assigns goal positions based on robot IDs, and (b) a minimal conflict strategy that assigns goals to minimize trajectory intersections, assuming straight-line movement from start to goal.
  • Motion Generation: Generates feasible paths from start to goal, considering environmental constraints. For each waypoint, the function not only outputs the position but also constraints such as maximum speed for approach.
This modular design allows the LLM to compose robust execution code for all selected action nodes, promoting reusability, flexibility, and scalability in multi-robot systems.
Trigger Condition Code Generation: For each action node, LLMs are prompted to implement the corresponding trigger condition. If no trigger is specified, the function always returns False. Otherwise, the LLM encodes the trigger logic using available environmental information, including robot and object states and region properties.
Action Node Finish Condition Code Generation: For each selected action node, LLMs generate code to evaluate completion based on its specified finish condition. For example, the task of an action node is to visit three points. The action node finishes when all three points are visited.
Mission Finish Condition Code Generation: Mission finish condition generation is similar in nature to composite task finish condition generation. However, in this case, we only need to prompt LLMs to generate the mission finish condition code once for the entire mission.
IV-C Execution
Once all code scripts are generated, the system begins or resumes execution. During execution, if the trigger condition is satisfied, we prompt LLMs to update the dependency analysis by removing the current task that caused the trigger and adding new tasks associated with this trigger. The updated dependency analysis, combined with the mission description, is then used to prompt LLMs to generate a new behavior tree. Afterward, LLMs select new ready-to-execute action nodes, and generate code for these nodes. If the finish condition of any action nodes is satisfied, LLMs update the behavior tree, including their parent nodes if the finish condition of these parent nodes solely relies on the completion of their children. LLMs then select new ready-to-run action nodes and generate corresponding code. If the finish condition of any composite nodes is satisfied, LLMs update the behavior tree by marking all its child nodes that have not yet started or are still running as Failure and marking the composite nodes as Success. The execution continues iteratively until the mission finish condition is fully satisfied.
Figure 3: Quantitative results for nine multi-robot missions. Missions #1–3 require LLMs to analyze simple missions and utilize the existing meta-behavior library for code generation. For example, mission #1 asks three robots on the left to visit the top left region, and three robots on the right to visit the three right regions separately. After these two tasks are completed, all robots then enclose the top left region. Missions #4-6 go further, requiring LLMs to synthesize code beyond the provided library (e.g., generating spiral trajectories). For example, mission #4 asks all robots to form an ’H’ first, followed by forming an ’I’. Since the library only supports point visits, LLMs need to compute each robot’s position. Missions #7-9 are the most complex, involving intricate dependencies, triggers, or adaptive behaviors. For example, mission #7 starts with asking the robots on the left to visit points next to the top green region and the three robots on the right to enclose the green region on the right. If any robot on the left comes within 1.5 units of the bottom blue region, a trigger redirects all three to the top-left region. Please check the website for more details.
V Experiments
The experimental study is designed to rigorously assess LAN2CB under a variety of multi-step, long-horizon missions. We focus on three questions.
  1. 1)
    How well does LAN2CB perform across diverse missions?
  2. 2)
    How does the use of a standardized template for mission descriptions affect the performance of LAN2CB compared to using raw natural language?
  3. 3)
    Can LAN2CB be reliably deployed on real robots?
V-A Performance Across Diverse Missions
To the best of our knowledge, no standard dataset exists for evaluating multi-step, long-horizon multi-robot mission execution under natural language instructions. Therefore, we curate a comprehensive mission suite comprising three groups (nine missions total), each designed to test the distinct capabilities of our system.
Mission Dataset: We organize the missions by the complexity of mission analysis and required code generation:
  • Category A: Basic Missions. These missions comprise multiple atomic tasks executed in a straightforward order, with no trigger conditions. Mission analysis and code generation are direct: all necessary information can be extracted from the natural language descriptions, enabling the LLM to generate input parameters for behavior library functions without additional reasoning. Examples: Assigning robots to visit predefined points. (See Missions #1, #2, and #3 in Fig. 3.)
  • Category B: Basic Missions w/ Advanced Code Generation. While maintaining a simple overall structure, these missions require advanced code generation. Certain behaviors are not directly supported by the behavior libraries, necessitating that the LLM synthesizes additional logic or routines. Example: Generating a spiral trajectory for robots when only straight-line motion is natively supported. (See Missions #4, #5, and #6 in Fig. 3.)
  • Category C: Advanced Missions. These missions demand advanced analysis and complex code generation. They may involve intricate dependencies, trigger conditions, or non-trivial mission-level finish criteria, such as region coverage requirements, dynamic triggers, or adaptive behaviors. (See Missions #7, #8, and #9 in Fig. 3.)
Evaluation Metrics: We quantitatively evaluate missions using the metrics: (1) Avg. Success Rate: The proportion of missions completed successfully as intended, without human intervention. (2) Avg. Token Number and Cost: The total number of LLM tokens used per mission and related cost. (3) Avg. Tokens per Error: The average number of tokens consumed per error given the output tokens of LLMs.
TABLE II: Quantitative results.
Mission Category W/ Template Avg. Success Rate Avg. Tokens ($10^{3}$) Avg. Cost (\$) Avg. Tokens / Err ($10^{3}$)
Input Output
Category A 0.87 48.70 12.33 0.20 92.48
1.00 52.83 14.53 0.22 -
Category B 0.73 40.53 9.97 0.16 37.39
0.93 35.25 8.54 0.14 128.10
Category C 0.73 44.24 10.06 0.17 37.73
0.87 58.12 14.86 0.24 111.45
Overall 0.78 44.49 10.79 0.18 55.87
0.93 48.73 12.64 0.20 119.78
*Avg. Tokens/Error is based on output tokens.
Results: We evaluate all missions using GPT-4.1, conducting each mission five times. Detailed results are provided in Tab. II. By standardizing raw mission descriptions from users with an expert-provided template, LAN2CB achieves a 100% success rate on Category A (basic missions), a 93% success rate on Category B (basic missions with advanced code generation) and 87% success rate on Category C (advanced missions). In addition, the system demonstrates strong robustness, achieving approximately 0.12 million tokens per error across all missions.
These results demonstrate that LAN2CB reliably analyzes mission requirements, including task sequencing, extraction of essential information from natural language instructions, selection of ready-to-run nodes, and dynamic updates to behavior tree-style representations. Notably, the high success rate in Category B and Category C highlights its ability to perform advanced code generation tasks. These include synthesizing robot trajectories not natively supported by the behavior library, generating custom formations, creating complex execution conditions such as specifying approach directions or target-following durations, generating triggers, and generating mission finish conditions. Overall, our results indicate that LAN2CB is both robust and adaptable for diverse multi-robot mission execution scenarios.
Figure 4: Snapshots of real-world experiments.   task not activated.   task running.   task finished.
V-B Impact of standardized template
To quantify the impact of using a template to standardize the mission descriptions on LAN2CB, we ablated the template-reformatting stage that aligns user-provided natural language mission descriptions to our predefined mission schema. As reported in Tab. II, this omission results in an average drop of approximately 15% in success rates and causes the error rate to become 2.14 times more frequent across all mission categories. These results highlight the critical role of structured prompting in maintaining robust multi-robot performance.
V-C Real World Demonstrations
We validated our framework through five real-world demonstrations using a team of ground robots across two robot platforms (see Fig. 4).
VI Conclusions
We presented an LLM-based framework, Language to Collective Behavior (LAN2CB), a pipeline that converts natural language into executable Python code for robot teams. Our experiments suggest that LAN2CB enables effective multi-robot coordination and remains flexible and extensible, even for multi-step and long-horizon missions. Additionally, we designed a dataset of natural language mission specifications for multi-robot coordination.
For future work, we aim to improve mission analysis through retrieval-augmented generation (RAG) [50], enabling LLMs to leverage past mission data more effectively. Furthermore, we plan to expand the meta-behavior library to support a broader and more diverse range of capabilities.
VII Acknowledgment
We thank Jiazhen Liu for assistance with the physical deployment.